At a dairy farm milk flows through a 22mm diameter pipe at an average velocity of 0.25m/s

p=1028kg/m^3 , u=3.0*10^-3 Pa.s

1) what is the volumetric flow rate in L/s

2) Determine whether the flow in the pipe is laminar or turbulent.support your answer.

3)Futhure along the pipe , the pipe dismeter reduces to 15mm. For the reduced section,determine whether the flow regime has remained the same as in (2) or has changed . Support your answers


1
Expert's answer
2021-06-08T10:21:02-0400

1)The volumetric flow rate in L/s

"V= \frac{AD^2}{32 \mu} \frac{\triangle P}{L}=\frac{ \pi r^2 D^2}{32} \frac{\triangle P}{1}"

"V=\frac{ \pi*(\frac{22}{2000})^2(\frac{22}{1000})^2}{32} \frac{3.0*10^{-3}}{1}=17.25*10^{-12}L/s"

2) Determine whether the flow in the pipe is laminar or turbulent.support your answer.

The Reynolds number determines whether the flow in a pipe is laminar or turbulent.

"Re=\frac{\rho u L}{\mu}=\frac{1028* 0.25 *1}{1}=257 < 4000" . laminar

3) "V=\frac{ \pi*(\frac{15}{2000})^2(\frac{15}{1000})^2}{32} \frac{3.0*10^{-3}}{1}=3.73*10^{-12}L/s"

The flow changed